OpenStudy (moonlight93):

non-right triangle Solve the missing sides and angles.. please help angle A= 30, side a = 3, and side c=6?

OpenStudy (anonymous):

 now use trig ratios to find the other side

OpenStudy (moonlight93):

I have to use law of sines and cosines to find missing angles and sides.... for one of the angles I got a 90 degree, it should be a non-right triangle.. I was hoping to see why I got that angle..... can some one please help me out... I need to know because I have a quiz over this

OpenStudy (anonymous):

In an non right triangle, there are six unknowns: the three angle measures and the three side lengths. To solve an oblique triangle you need one of the following sets of information: Two sides and an angle opposite one of the known sides Two angles and any side Two sides and their included angle All three sides If you know either (1) or (2), you can use the law of sines to solve the triangle. If you know (3) or (4), you must tag-team with the law of cosines and then the law of sines to find the solution.

OpenStudy (anonymous):

for Laws of sine and cosine read this http://www.sparknotes.com/testprep/books/sat2/math2c/chapter9section9.rhtml

OpenStudy (moonlight93):

angle C= 90, Angle B= 60 ??

OpenStudy (anonymous):

Law of sine \[a/ \sin (A) = c / \sin (C)\] \[3/\sin (30) = 5/\sin(C)\] \[3/(1/2) = 5/\sin(C)\] \[5/6 = \sin(C)\] Noe you can find \[\sin^{-1} (5/6)\]

OpenStudy (moonlight93):

how about the sides?

OpenStudy (moonlight93):

by cosin law?

OpenStudy (moonlight93):

side b= 6.87

OpenStudy (moonlight93):

RIGHT?

OpenStudy (moonlight93):

C=6 NOT 5

OpenStudy (anonymous):

Sure??? Cause then \[6= 6/\sin(C)\] \[C= \sin^{-1} (6/6)\] \[C= \sin^{-1} (1)\] then C = 90 degree

OpenStudy (moonlight93):

could it be 90 degree and the whole function has to be non right trianfle

OpenStudy (triciaal):